# Distributionally Robust Policy Learning under Concept Drifts

### v1

@@ Appendix C Proof of technical lemmas
We have thus completed the proof part (1) of Lemma B.1.
@@
where $\tilde{\bm{\theta}}=t\bm{\theta}(x)+(1-t)\bm{\theta}_{\pi}^{*}$ for some $t\in[0,1]$. Let $\xi_{1}=\min(\underline{\alpha},\bar{\eta})/2$. When $\|\bm{\theta}-\bm{\theta}_{\pi}^{*}\|_{L_{\infty}}\leq\xi_{1}$, we have $|\tilde{\alpha}-\bm{\alpha}^{*}_{\pi}(x)|\leq\xi_{1}$ and $|\tilde{\eta}-\bm{\eta}^{*}_{\pi}(x)|\leq\xi_{1}$. Plugging the expressions of the gradient in Equation (C.1), we have
@@
Using the sequential approximation operators, we decompose the inner expectation term in (B.3.2) (Rademacher complexity) as
@@
To proceed, we shall use the following lemma, whose proof is deferred to Appendix C.3.
Lemma C.2.
@@
C.4 Proof of Lemma B.3
By Yang et al., (2022, Lemma B12), $g_{\delta}(q)$ is differentiable in $q$, and
@@
Also by Yang et al., (2022, Lemma B12), $g_{\delta}(q)$ is convex in $q$, so $g_{\delta}^{\prime}(q)$ is increasing in $q$. Since $q\in[0.4,0.6]$, $g^{\prime}_{\delta}(q)\geq g^{\prime}_{\delta}(0.4)$. From the dual form, we can check that $g(0.4)\geq 0.1$. Plugging in $q=0.4$, we have

### v2

@@ Appendix F Proof of technical lemmas
We have thus completed the proof part (1) of Lemma D.1.
@@
where $\tilde{\bm{\theta}}=t\bm{\theta}(x)+(1-t)\bm{\theta}_{\pi}^{*}$ for some $t\in[0,1]$. Let $\xi_{1}=\min(\underline{\alpha},\bar{\eta})/2$. When $\|\bm{\theta}-\bm{\theta}_{\pi}^{*}\|_{L_{\infty}}\leq\xi_{1}$, we have $|\tilde{\alpha}-\bm{\alpha}^{*}_{\pi}(x)|\leq\xi_{1}$ and $|\tilde{\eta}-\bm{\eta}^{*}_{\pi}(x)|\leq\xi_{1}$. Plugging the expressions of the gradient in Equation (F.1), we have
@@
Using the sequential approximation operators, we decompose the inner expectation term in (D.3.2) (Rademacher complexity) as
@@
To proceed, we shall use the following lemma, whose proof is deferred to Appendix F.3.
Lemma F.2.
@@
F.4 Proof of Lemma D.3
By Yang et al. (2022, Lemma B12), $g_{\delta}(q)$ is differentiable in $q$, and
@@
Also by Yang et al. (2022, Lemma B12), $g_{\delta}(q)$ is convex in $q$, so $g_{\delta}^{\prime}(q)$ is increasing in $q$. Since $q\in[0.4,0.6]$, $g^{\prime}_{\delta}(q)\geq g^{\prime}_{\delta}(0.4)$. From the dual form, we can check that $g(0.4)\geq 0.1$. Plugging in $q=0.4$, we have